OpenStudy (anonymous):

Solve. A money market fund advertises a simple interest rate of 6%. Find the total amount received on an investment of $15,000 for 3 months. A. $17,700.00 B. $17,250.00 C. $225.00 D. $15,225.00

OpenStudy (anonymous):

Simple interest is the amount, times the percent interest for one time period, times the number of periods. So 15,000 times 3 months times 0.06=2,700. So 15,000+2,700= 17,700.

OpenStudy (anonymous):

The answer should be A by the way.

OpenStudy (anonymous):

actually its d.

OpenStudy (lilg132):

how?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

the answer said so lol

OpenStudy (lilg132):

your answers wrong then lol

OpenStudy (anonymous):

Do they show how?

OpenStudy (anonymous):

no it from an automatic test.

OpenStudy (lilg132):

hmmm

OpenStudy (anonymous):

If not, this is a question you need to ask your teacher. Scantrons mess up too.

OpenStudy (anonymous):

you dont times 15000 from 3 your times 15000*.06 then (3/12)

OpenStudy (anonymous):

its not on a scan tron its online

OpenStudy (anonymous):

ohhhhh

OpenStudy (anonymous):

They have to tell you that...

OpenStudy (lilg132):

ye so you mean you then times it by 4?

OpenStudy (anonymous):

yes i think so

OpenStudy (lilg132):

18600

OpenStudy (lilg132):

none of the options

OpenStudy (anonymous):

its d. trust me. Ill ask my professor tomorrow

OpenStudy (lilg132):

i got it!

OpenStudy (anonymous):

Simple interest has a definition of being time periods under a year. "Simple interest is normally used for a single period of less than a year, such as 30 or 60 days." http://www.getobjects.com/Components/Finance/TVM/iy.html So I assumed that you don't have to take it out of a year. Meh. Learned something.

OpenStudy (lilg132):

15,000 * 0.06 = 900 900 / 12 = 75 75 * 3 = 225 15,000 + 225 = 15,225

OpenStudy (lilg132):

sorry about that

OpenStudy (anonymous):

omg thats awesome i forgot you can do 365

OpenStudy (anonymous):

omg you rock lilg

OpenStudy (anonymous):

and for months 12

OpenStudy (lilg132):

:p